OpenStudy (anonymous):

y varies inversely as twice x. When x = 9, y = 2. Find y when x =

jimthompson5910 (jim_thompson5910):

"y varies inversely as twice x" means that y = k/(2x) y = k/(2x) 2 = k/(2(9)) 2 = k/18 2*18 = k 36 = k k = 36 --------------------- y = k/(2x) y = 36/(2x) y = 18/x So the equation is y = 18/x